OpenStudy (kenda):

the graph of a relation defined by 3y^2=x

OpenStudy (jdoe0001):

have you done quadratics yet?

OpenStudy (kenda):

yes

OpenStudy (anonymous):



OpenStudy (jdoe0001):

ok... well, this one is just a parabola... and assuming you know the graph of a parabola... so, say \(\bf y= \pm x^2\) is a vertical parabola, going up and \(\bf x=\pm y^2\) is a horizontal parabola, going sideways what you have is a horizontal parabola... in this case.. \(\large 3y^2=x\implies 3(y-0)^2=(x-0)\\\qquad \begin{array}{llll} &&\textit{horizontal shrink}\\ &&\downarrow \\ (x-&0)=&3(y-&0)^2\\ center(&\uparrow, &&\uparrow ) \end{array} \)